40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17004     Entry time: Thu Jul 14 19:56:15 2022
Author: rana 
Type: Update 
Category: IOO 
Subject: mc wfs demod 

It looks like Tomislav's measurements of the WFS demod board noise were actually of the cable that goes from the whitening to the ADC. So the huge low frequency excess that he saw is not due to wind, but just the inverse whitening of the digital system?

In any case, today, I looked at the connections from the Whitening to the ADC. It goes through an interface chassis to go from ribbon to SCSI. The D-Sub connectors there have the common problem in many of the LIGO D-sub connectors: namely that the strain relief nuts are too tall and so the D connector doesn't seat firmly - its always about to fall out. JC, can you please take a look at this and order a set of low profile nuts so that we can rework this chassis? Its the one between the WFS whitening and the SCSI cables which go to the ADCs.

After pushing them in, I confirmed that the WFS are working, by moving all 6 DoF of the MC mirrors via bias slider, and looking at the step responses (attached). As you can see, all sensors see all mirrors, even if they are noisy.

Next up: get a breakout for the demod output connector and measure the noise there.

For today, I aligned the IMC by hand, then centred the WFS beams by unlocking the IMC and aligning the bright beam. I noticed that the WFS1 beam was being dumped randomly, so I angled the WFS1 by ~3 deg and dumped the specular reflection on a razor blade dump. To handle the sign change in the MC1 actuation (?), I changed the sign in the MC1 ASC filter banks. MCWFS loops still nto closing, but they respond to mirror alignment.

Attachment 1: mcwfs-steps.pdf  16 kB  Uploaded Thu Jul 14 21:14:23 2022  | Hide | Hide all
mcwfs-steps.pdf
ELOG V3.1.3-